# 3VL: Using Trees to Improve Vision-Language Models’ Interpretability

### v1

@@ V-J Limitations
Note that although 3VL improves upon vanilla CLIP in all words, we sometimes see an increase in bias towards one word of a pair: Some pairs of words have a big gap between the fail rate when one of the words is the positive one and the other one is the negative word, compared to the fail rate with the same pair but with the positive and negative words switched. For example, as shown in Table VIII, when the word “small” is positive and the word “large” is negative CLIP has a 50% fail rate, but when the word “large” is the positive one and the word “small” is negative the fail rate is just above 25% which suggests a very strong bias of CLIP to the word “large” over “small”. With 3VL these fail rates are reduced to 35.8% and 22.1% and the gap is smaller. Yet, with the pairs “on”-“off” and “young”-“old” the bias is reduced a lot less with 3VL. Moreover, for the pairs “down”-“up”, “black”-“white”, and “man”-“woman” the gap even increases, although the model makes less mistakes overall. This is because the COCO train set contains an inherent large bias to one word from each of these pairs.
Having said that, we believe that the tree structure used with 3VL can be employed in order to decrease this gap. Given that we have the tree structure, we encourage a greater penalty on the tree leaves for text that appears less in the training data. We believe that our proposed tree augmentation method can be utilized to reduce biases in VLM training but we leave this direction to future work.
@@ VI Conclusion
In this work, we propose a novel approach to address the limitations of VLMs in Compositional Language Concepts understanding and to enhance the interpretability of VLMs by leveraging the hierarchical structure of natural language. Our approach offers a richer exploration of the text space using several levels of incremental text augmentations from coarse to fine-grained, and in the process allows for better interpretation and analysis of model decisions. We complement our method with image heatmap relevancy maps of positive and negative texts and use this to identify further weaknesses of VLMs. We propose a new inference technique that effectively filters out the noise and irrelevant information from the input, allowing us to gain insights into the underlying failure modes of the model and achieve state-of-the-art performance on popular CLC benchmarks. We believe that our proposed tree augmentation method can be utilized to reduce biases in VLM training but we leave this direction to future work. Our work has important implications for the development of more interpretable and effective VLMs for natural language understanding and computer vision. Note that our proposed solution is generic. We have demonstrated it here with CLIP. Yet, a future work can easily incorporate it in other VLMs.

### v2

@@ V-J Limitations
Note that although 3VL improves upon vanilla CLIP in all words, we sometimes see an increase in bias towards one word of a pair: Some pairs of words have a big gap between the fail rate when one of the words is the positive one and the other one is the negative word, compared to the fail rate with the same pair but with the positive and negative words switched. For example, as shown in Table IX, when the word “small” is positive and the word “large” is negative CLIP has a 50% fail rate, but when the word “large” is positive and the word “small” is negative the fail rate is just above 25% which suggests a very strong bias of CLIP to the word “large” over “small”. With 3VL these fail rates are reduced to 35.8% and 22.1% and the gap is smaller. Yet, with the pairs “on”-“off” and “young”-“old” the bias is reduced a lot less with 3VL. Moreover, for the pairs “down”-“up”, “black”-“white”, and “man”-“woman” the gap even increases, although the model makes fewer mistakes overall. This is because the COCO train set contains an inherent large bias to one word from each of these pairs.
Having said that, we believe that the tree structure used with 3VL can be employed to decrease this gap. Given that we have the tree structure, we encourage a greater penalty on the tree leaves for text that appears less in the training data. We believe that our proposed tree augmentation method can be utilized to reduce biases in VLM training but we leave this direction to future work.
@@ VI Conclusion
In this work, we propose a novel approach to address the limitations of VLMs in Compositional Language Concepts understanding and to enhance the interpretability of VLMs by leveraging the hierarchical structure of natural language. Our approach offers a richer exploration of the text space using several levels of incremental text augmentations from coarse to fine-grained, and in the process allows for better interpretation and analysis of model decisions. We complement our method with image heatmap relevancy maps of positive and negative texts and use this to identify further weaknesses of VLMs. We propose a new inference technique that effectively filters out the noise and irrelevant information from the input, allowing us to gain insights into the underlying failure modes of the model and achieve state-of-the-art performance on popular CLC benchmarks. We believe that our proposed tree augmentation method can be utilized to reduce biases in VLM training but we leave this direction to future work. Our work has important implications for the development of more interpretable and effective VLMs for natural language understanding and computer vision. Note that our proposed solution is generic. We have demonstrated it here with CLIP. Yet, future work can easily incorporate it in other VLMs.
Appendix A User Study Screenshots
In Fig 10 we provide several screenshots from our user study. They demonstrate the way the images with the questions were presented to the users in order to measure the interpretability of the different approaches.
[Uncaptioned image] Nir Yellinek is a M.Sc graduate from Tel-Aviv University. He received his B.Sc (2011) and M.Sc. (supervision by Prof. Raja Giryes, 2023) degrees from the school of electrical engineering at Tel-Aviv University. His research interests include deep learning, self-supervised, multi-modal and AI alignment.
[Uncaptioned image] Leonid Karlinsky is a Principal Research Scientist (STSM) in the MIT-IBM lab. Prior to that Leonid led the AI Vision research group in the Multimedia department @ IBM Research AI. Leonid joined IBM Research in July 2015. Before joining IBM, he served as a research scientist in Applied Materials, Elbit, and FDNA. He is actively publishing, reviewing, and performing occasional chair duties at ECCV, ICCV, CVPR, ICLR, AAAI, WACV, and NeurIPS, and is serving as an IMVC steering committee member for the past 6 years. During his time at IBM, Leonid has co-authored over 25 research papers on his research is in the areas of augmented reality, medical applications, self-supervised, cross-domain, multi-modal, and few-shot learning. He received his PhD degree at the Weizmann Institute of Science, supervised by Prof. Shimon Ullman.
[Uncaptioned image] Raja Giryes is an associate professor in the school of electrical engineering at Tel Aviv University. His research interests lie at the intersection between deep learning and signal processing. Raja received the EURASIP best P.hD. award, the ERC-StG grant, Maof young faculty fellowship (2016-2019), VATAT postdoctoral scholarship (2014-2015), the TI Excellence in Signal Processing Award (ESPA)(2008), the Azrieli Fellowship (2010-2013). He is IEEE Senior Member and member of the Israeli young academy. He has 150+ publications in top venues with 9500+ citations. 30 MSc and 6 PhD students already graduated from his group. Two of them hold faculty positions at Bar Ilan University and University of Chicago. He is an associate editor in IEEE TPAMI, IEEE TIP and Elsevier Pattern Recognition and an area chair in NeurIPS. He organized workshops and tutorials on various aspects of deep learning both internationally and locally including in ICML, CVPR, ECCV and ICCV